# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio Choice

### v5

@@ Abstract
Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality. Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a few state variables ($\sim$3-6 limit in prior studies). To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework. Our framework accurately captures both myopic demand and complex intertemporal hedging demands, a feat often elusive for other methods in high-dimensional settings. P-PGDPO delivers near-optimal policies, offering a practical and powerful alternative for a broad class of high-dimensional continuous-time control problems. PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to directly inform neural network-based policy learning. A key contribution is our highly efficient Projected PG-DPO (P-PGDPO) variant. This approach uniquely utilizes BPTT to obtain rapidly stabilizing estimates of the Pontryagin costates and their crucial derivatives with respect to the state variables. These estimates are then analytically projected onto the manifold of optimal controls dictated by PMP’s first-order conditions, significantly reducing training overhead and enhancing accuracy. This enables a breakthrough in scalability: numerical experiments demonstrate that P-PGDPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables).
@@ 1 Introduction
The theory of dynamic consumption and portfolio choice has been a cornerstone of financial economics, providing fundamental insights into asset pricing and long-term investment behavior (Merton, , 1973; Campbell and Viceira, , 1999, 2001). Despite its theoretical elegance and contributions to our understanding, the theory faces a major obstacle in practical implementation: the lack of scalable methods for solving high-dimensional portfolio choice problems involving many assets and state variables.
The Bellman equation associated with such problems is subject to the curse of dimensionality, and obtaining solutions in realistic settings is often infeasible. Existing solution methods are typically complex, ad hoc, and do not scale well with dimensionality. This is one reason why Markowitz’s static mean-variance framework remains dominant in practice, despite its limitations and the greater realism offered by dynamic models. [^1] Cochrane wrote “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice. ” (Cochrane, , 2022, p.4)
In this paper, we make a technological contribution by proposing a solution approach that leverages machine learning with neural networks to solve high-dimensional dynamic portfolio choice problems, where both the number of assets $n$ and the number of state variables $k$ are large. There are two primary challenges inherent in such problems. The first is the well-known curse of dimensionality: computational methods often fail to scale efficiently as the dimensionality of the problem increases in $n$ and $k$. The second, and more subtle, challenge arises from the specific structure of dynamic portfolio choice itself. Unlike other applications of reinforcement learning—such as Atari, Go, or autonomous driving—where achieving a near-optimal policy is often sufficient for strong empirical performance, portfolio optimization requires much greater precision. A well-established theoretical benchmark exists, dating back to the seminal work of Merton, (1973), which demonstrates that the optimal portfolio can be decomposed into two critical components: the myopic demand and the intertemporal hedging demand. The myopic component represents the demand for a mean-variance efficient portfolio given current investment opportunities, while the intertemporal component hedges against expected future changes in those opportunities driven by evolving economic conditions. Accurately identifying both components is essential for evaluating investor behavior and designing long-term investment strategies—yet it remains a formidable challenge, particularly in high-dimensional settings. Our proposed approach is specifically designed to address both challenges: overcoming the curse of dimensionality and enabling the accurate decomposition of optimal portfolios into their myopic and intertemporal components.
We adopt a continuous-time framework in which asset return means and covariances are driven by a $k$-dimensional state vector, with the state variables evolving as Itô diffusions, following Merton, (1973). This modeling approach benefits from the rich toolkit of stochastic calculus—tools generally unavailable in discrete-time formulations. [^2] See Duarte et al., (2024) for a discussion of the advantages of continuous-time models over discrete-time models. In the continuous-time setting, (quasi-)analytic solutions are available under affine dynamics, where the Hamilton–Jacobi–Bellman (HJB) equation reduces to a system of ordinary differential equations (Kim and Omberg, , 1996; Liu, , 2007; Buraschi et al., , 2010). However, for general non-affine dynamics, no universal solution method exists. To address this limitation, we introduce a machine learning–based approach that applies to arbitrary market dynamics. Neural networks, known for their ability to approximate complex high-dimensional functions, are used to parameterize the policy directly, enabling us to capture intricate dependencies on both wealth and the exogenous state variables—even when the dimension $k$ is large.
While we focus on Brownian motion as the source of return risk, the methods can be extended to accommodate jumps. Furthermore, if a discrete-time model is viewed as an approximation of a continuous-time model, our approach can be adapted accordingly.
Our approach leverages Pontryagin’s Maximum Principle (PMP) as its theoretical foundation, in contrast to the more commonly used dynamic programming (DP) framework based on the Hamilton–Jacobi–Bellman (HJB) equation. Building on PMP, we propose two methods: Pontryagin-Guided Direct Policy Optimization (PGDPO) and its variant, Projected PGDPO (P-PGDPO).
@@
PMP characterizes optimal controls through a coupled system of forward and backward differential equations. The forward equation governs the evolution of wealth and state variables driving asset returns, which we implement via Monte Carlo simulation. This allows for direct computation of the expected utility and avoids the need to estimate the value function across future periods—a core but computationally intensive requirement of DP-based approaches.
@@
The backward equation governs the dynamics of the costates (or adjoint variables), which correspond to the first-order partial derivatives of the value function with respect to the state variables. Modern machine learning frameworks such as PyTorch allow us to compute these costates and their derivatives, the second-order partial derivatives of the value function, automatically through backpropagation through time (BPTT) and automatic differentiation. We demonstrate that a single forward simulation pass, coupled with BPTT, yields unbiased estimates of the costate processes required by PMP. As a result, our approach avoids the need for explicit solution of the backward equation or approximation of the value function, thereby sidestepping major bottlenecks of DP. Costates—as well as the value function—emerge as natural byproducts of the learning process.
In PGDPO, the control policy—comprising consumption and portfolio allocations—is parameterized by neural networks. The expected utility, defined as the discounted sum of utility from consumption and terminal wealth, is then directly maximized using stochastic gradient ascent, where gradients are computed using BPTT. A key computational advantage of our approach lies in its complete avoidance of HJB residual minimization, where the residual is defined as the cumulative squared error of the HJB equation. Computing the HJB residual typically requires evaluating first- and second-order partial derivatives of the value function at each time–state grid point and aggregating the errors over time, incurring high computational cost. Duarte et al., (2024) mitigate this by replacing full gradient computation with the second derivative of a carefully constructed univariate function. In contrast, our method bypasses HJB residual computation entirely, resulting in significant computational savings.
Conceptually, our approach shares similarities with recent methods that parameterize control policies via neural networks and directly optimize the expected objective function through simulation and gradient ascent, thereby circumventing the need to solve the HJB equation (Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024, e.g., for portfolio problems with transaction costs or Heston dynamics). However, our PMP-based PGDPO framework distinguishes itself by being explicitly guided by continuous-time optimality conditions derived from PMP.
Moreover, unlike standard reinforcement learning methods such as Proximal Policy Optimization (PPO) or Trust Region Policy Optimization (TRPO), which rely on sampled estimates of policy gradients, our method computes exact gradients via BPTT. This yields lower-variance updates and enhances computational efficiency, particularly in high-dimensional settings (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
The second method, P-PGDPO, builds on PGDPO by substituting the estimated derivatives of the value function directly into the structural first-order conditions from PMP. These conditions take the form derived originally by Merton, (1973), explicitly decomposing the optimal portfolio into myopic and intertemporal hedging demands. We demonstrate the numerical accuracy of our methods by comparing the solutions they produce to a quasi-analytic benchmark under affine dynamics, where the true solution can be obtained by solving a system of ordinary differential equations.
The two components of the optimal portfolio—myopic and intertemporal hedging demands—are not merely of theoretical interest; their accurate identification requires high-fidelity solutions. As shown by Constantinides, (1986), transaction costs generate only second-order utility losses, allowing portfolios to deviate substantially from the frictionless optimum while still appearing nearly optimal in terms of utility. Similarly, Cochrane, (1988) demonstrates that first-order deviations from the optimal policy can result in only second-order losses in utility.
This disconnect between utility value and policy structure poses a major challenge for traditional reinforcement learning algorithms. These methods often converge to policies that are near-optimal in utility but structurally far from the true optimum—particularly in high-dimensional settings. In such cases, neural networks must search for optimal policies over an $n$-dimensional control space, where $n$ denotes the number of risky assets. The resulting exponential growth in complexity renders standard policy optimization techniques computationally infeasible as $n$ increases. Our numerical experiments confirm that increasing the number of risky assets $n$ presents a significantly greater challenge than increasing the number of state variables $k$ in high-dimensional settings.
The P-PGDPO method is motivated by the empirical observation that value function derivatives—i.e., costates—are typically easier to learn than the policy functions themselves. In practice, these partial derivatives tend to stabilize much earlier during training. Once the costate estimates have stabilized—usually after a short warm-up phase—the method can bypass the potentially suboptimal outputs of the policy networks. Instead, the costates and their derivatives are directly substituted into the first-order conditions (FOCs) from PMP, whose structural form closely mirrors the conditions originally derived by Merton, (1973). This effectively projects the learned costate information onto the manifold of controls that satisfy PMP’s necessary conditions. The resulting control policy adheres closely to the structural form of the true optimal solution dictated by PMP. It is recovered quasi-analytically at each time–state point, requiring minimal additional training and incurring substantially reduced computational cost. Our PMP-based PGDPO framework thus leverages BPTT-derived estimates of costates and their derivatives, enhancing scalability while providing a principled mechanism for enforcing first-order optimality conditions.
@@
We provide a theoretical justification for the P-PGDPO method in Theorem 1. When the objective function surface is relatively flat near the optimum, the first-order condition (FOC) violation for the policy obtained in the warm-up stage tends to be small. The theorem states that if this FOC deviation is small and the costate estimates obtained via BPTT are sufficiently accurate, then the control policy recovered in the second stage is close to the true optimal policy.
We demonstrate the accuracy of our numerical methods by solving a problem in which the state process follows a multi-factor Ornstein–Uhlenbeck (OU) process—a special case of affine dynamics—for which quasi-analytic benchmark solutions can be obtained by solving a system of ordinary differential equations (see Section 2.3.2). We also compare our methods with the Deep Backward Stochastic Differential Equation (Deep BSDE) approach, which is representative of existing techniques for solving high-dimensional partial differential equations, particularly the Hamilton–Jacobi–Bellman (HJB) equations arising in DP. In addition, we apply our second method, P-PGDPO, to a non-affine setting by adding non-linear terms to the expected returns of risky assets, thereby demonstrating its applicability to a broader class of dynamics.
Our numerical experiments (Section 4) show that our first method, PGDPO, produces solutions comparable to those of the Deep BSDE method in terms of mean-squared errors. In contrast, the second method, P-PGDPO, yields significantly smaller errors. This improvement arises because PGDPO inherits the common limitation of policy optimization algorithms—difficulty in discovering the true optimum—whereas P-PGDPO overcomes this issue by directly exploiting the structural form of the first-order conditions from Pontryagin’s Maximum Principle (PMP).
P-PGDPO not only scales to problems involving dozens of assets and state variables (e.g., $n=50$, $k=10$), but also accurately disentangles and reconstructs complex intertemporal hedging demands. Furthermore, our results demonstrate that the method generates stable solutions in non-affine settings and converges smoothly to the benchmark affine solution as the coefficients of the nonlinear terms approach zero.
By contrast, the Deep BSDE approach, while successful in other high-dimensional applications such as option pricing and hedging, fails to capture the full structure of the optimal policy in our setting—particularly the intertemporal hedging components.
The method’s superior performance—both in scalability and in fidelity to the full structure of the optimal policy—highlights the transformative potential of leveraging PMP through BPTT-derived costate estimates for a broad class of high-dimensional control problems that admit stable costate representations.
There is a substantial literature on dynamic portfolio choice, aimed at obtaining concrete solutions either analytically or numerically. With the exception of the continuous-time models with affine dynamics discussed above, most existing contributions are formulated in discrete time. For example, Campbell and Viceira, (1999) and Campbell and Viceira, (2001) employ log-linear approximations to the dynamic budget constraint to obtain approximate closed-form solutions. Balduzzi and Lynch, (1999), Lynch and Balduzzi, (2000), and Lynch, (2001) apply discrete-space approximations to solve portfolio choice problems with transaction costs. Garlappi and Skoulakis, (2010) develop a DP approach based on state variable decomposition.
Brandt et al., (2005) propose a Monte Carlo simulation method for a discrete-time portfollio choice model that projects simulated values of relevant functions—such as the marginal value of wealth—onto a pre-specified set of basis functions. In this sense, their method shares similarities with the forward simulation and auto-differentiation steps in our PG-DPO framework. However, their approach does not explicitly leverage PMP or utilize the representational power of neural networks. As a result, it is not easily scalable and lacks the flexibility for generalization. In contrast, our method is PMP-guided, scalable, and extensible to broader classes of control problems.
In continuous time, there has been significant technological advancement in solving partial differential equations, which can be applied to the HJB equation arising from DP. Notable methods include Deep BSDE, Deep Galerkin, Physics-Informed Neural Networks (PINNs), Deep Splitting, and Deep Branching, among others (Han et al., , 2018; Sirignano and Spiliopoulos, , 2018; Raissi et al., , 2019; Beck et al., , 2021; Nguwi et al., , 2024).
@@
In parallel, substantial progress has been made in solving optimal control problems through the development of policy evaluation, policy gradient, $q$-learning, and PPO/TRPO methods in stochastic environments (Wang et al., , 2020; Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou, , 2023; Zhao et al., , 2023). Wang and Zhou, (2020); Dai et al., (2023); Huang et al., (2024) apply these techniques to dynamic mean-variance analysis, while Dai et al., (2025) use them to solve a Merton-type portfolio optimization problem. These approaches involve learning the asset return process from observed return data. Geng et al., (2023) propose the Factor Learning Portfolio Optimization (FaLPO) method for solving a Merton-type portfolio optimization problem with stochastic factors. Their approach combines policy learning with model calibration. By contrast, we assume that the data-generating process is known and do not address the problem of learning from data. Instead, our focus is on a more general consumption and portfolio choice problem involving many risky assets and multiple state variables, whereas their analysis is limited to a small number of risky assets and state variables.
Davey and Zheng, (2022) develop a BSDE-based method for portfolio optimization under portfolio constraints that is also based on the maximum principle, similar in spirit to our work. However, their method requires minimizing the terminal value error (as in Deep BSDE), maximizing the Hamiltonian at each state and time, and solving a BSDE system. In contrast, our method directly maximizes expected utility without penalizing approximation errors and without solving the backward equation explicitly, resulting in a significantly simpler and more efficient approach.
More recently, Duarte et al., (2024) propose a machine learning approach for infinite-horizon continuous-time problems using DP, while Cheridito et al., (2025) develop a method that combines generalized policy iteration with PINNs to solve finite-horizon continuous-time DP problems. By contrast, our method is based on PMP, rather than DP. The aforementioned DP-based methods avoid the explicit evaluation of first- and second-order partial derivatives of the value function by instead computing the second derivative of a univariate function when assessing the HJB residual. In contrast, our approach computes these partial derivatives directly—via automatic differentiation of costate variables—providing crucial information for accurately capturing intertemporal hedging components. This ability to compute the value function’s derivatives with high accuracy enables precise estimation of both the intertemporal hedging terms and the myopic component of the optimal portfolio. [^3] Cochrane emphasizes the critical role of partial derivatives in determining the optimal portfolio for long-term investors, particularly when discussing the implementation challenges of Merton’s dynamic portfolio theory—a point we partially quoted earlier: “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice… What are the partial derivatives of the investor’s value function that guide state-variable investments? That is harder still…” (Cochrane, , 2022, pp. 4–5). To our knowledge, the two aforementioned papers do not demonstrate accurate recovery of intertemporal hedging terms in settings with both large numbers of assets and state variables.
To our knowledge, the existing literature has demonstrated the ability to solve problems involving at most a dozen risky assets and a dozen state variables—with two notable exceptions. Davey and Zheng, (2022) consider a setting with 20 assets, but under a constant investment opportunity set, i.e., without state variables influencing asset returns. [^4] Our companion paper demonstrates that our method is applicable to constrained portfolio optimization problems involving up to 1000 assets (Huh et al., , 2025). Cheridito et al., (2025) address a case with 25 assets, though without providing evidence regarding the accuracy of the resulting solution. In contrast, our method substantially improves scalability, handling much larger values of $n$ and comparably high values of $k$, while preserving accuracy even in high-dimensional settings.
The main contributions of this paper can be summarized as follows:
  • We introduce Pontryagin-Guided Direct Policy Optimization (PG-DPO), and its highly efficient “Projected PG-DPO” (P-PGDPO) variant that uses PMP structure (specifically FOCs) and BPTT-derived costates to solve complex continuous-time control problems.
  • We demonstrate through numerical experiments that P-PGDPO not only breaks the dimensional barrier in portfolio optimization (up to n=50 assets, k=10 factors), but critically, accurately captures complex intertemporal hedging demands in these high-dimensional settings, a capability not typically demonstrated by alternative deep learning approaches.

### v6

@@ Abstract
We introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework for high-dimensional continuous-time portfolio choice. Our approach combines Pontryagin’s Maximum Principle (PMP) with backpropagation through time (BPTT) to directly inform neural network–based policy learning, enabling accurate recovery of both myopic and intertemporal hedging demands—an aspect often missed by existing methods. Building on this, we develop the Projected PG-DPO (P-PGDPO) variant, which achieves near-optimal policies with substantially improved efficiency. P-PGDPO leverages rapidly stabilizing costate estimates from BPTT and analytically projects them onto PMP’s first-order conditions, reducing training overhead while improving precision. Numerical experiments show that PG-DPO matches or exceeds the accuracy of Deep BSDE, while P-PGDPO delivers significantly higher precision and scalability. By explicitly incorporating time-to-maturity, our framework naturally applies to finite-horizon problems and captures horizon-dependent effects, with the long-horizon case emerging as a stationary special case.
Keywords: Multi-asset Portfolio choice, Pontryagin’s Maximum Principle, direct Policy Optimization, model-based reinforcement learning
[…]
[A-]Online-Appendix
@@ 1 Introduction
The continuous-time consumption and portfolio choice problem has been a cornerstone of financial economics, offering fundamental insights into asset pricing and long-term investment behavior (Merton, , 1973; Campbell and Viceira, , 1999, 2001). Despite its theoretical elegance, the practical implementation of this framework has remained exceptionally challenging. Two obstacles are particularly salient. First, portfolio choice problems involving many risky assets and state variables are subject to the curse of dimensionality (Bellman, , 1966). Traditional dynamic programming (DP) approaches, built on solving the Hamilton–Jacobi–Bellman (HJB) equation, readily become intractable as dimensionality increases. This is one reason why Markowitz’s static mean–variance framework still remains dominant in practice, despite its limitations and the greater realism offered by dynamic models. [^1] Cochrane wrote: “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice. (Cochrane, , 2022, p.4)” The appeal of the mean-variance model lies in its tractability, as it reduces portfolio choice to a single-period quadratic optimization problem that remains computationally feasible even with many assets. However, it ignores intertemporal hedging motives, relies on a restrictive single-period formulation, and assumes normally distributed returns so that only means and variances matter. Moreover, it disregards the dynamic evolution of investment opportunities and practical frictions such as transaction costs or portfolio constraints, and suffers from a lack of time consistency in multi-period settings. Consequently, most continuous-time models in the literature have been confined to highly stylized, low-dimensional settings. Second, even when solutions are available, they often fail to recover the structural decomposition of the optimal policy. A well-established theoretical benchmark, dating back to the seminal work of Merton, (1973), demonstrates that the optimal portfolio can be decomposed into two critical components: the myopic demand and the intertemporal hedging demand. The myopic component represents the demand for a mean-variance efficient portfolio given current investment opportunities, while the intertemporal component hedges against expected future changes in those opportunities driven by evolving economic conditions. Accurate identification of both components is crucial for evaluating investor behavior and designing long-term investment strategies.
Recent advances in machine learning have introduced new PDE-based solvers, including Physics-Informed Neural Networks (PINNs), the Deep BSDE method (Han et al., , 2018), and Deep Galerkin Networks (DGNs). While these approaches help mitigate the curse of dimensionality, they focus on approximating the value function and are not designed to recover the full structure of the optimal portfolio, particularly the intertemporal hedging demand. Moreover, contributions such as Duarte et al., (2024) restrict attention to infinite-horizon problems, where the optimal solution is stationary and time-independent. Although suitable for certain theoretical contexts, such methods are ill-suited for practical applications such as retirement planning, endowment management, or finite-maturity products, where horizon effects and time-varying hedging demands play a central role.
To address these limitations, we propose a new framework, Pontryagin-Guided Direct Policy Optimization (PG-DPO). Our approach integrates Pontryagin’s Maximum Principle (PMP) with policy networks trained via backpropagation through time (BPTT), enabling scalable and structurally accurate solutions to high-dimensional continuous-time portfolio choice problems.
@@
PMP characterizes optimal controls through a coupled system of forward and backward equations. The forward equation governs the evolution of wealth and state variables that drive asset returns, which we simulate using Monte Carlo methods. This allows direct computation of expected utility and removes the need to estimate the value function over future periods—a core but computationally intensive step in DP-based approaches. Accordingly, our framework relies solely on policy networks, without requiring a value network. The backward equation governs the dynamics of the costates (or adjoint variables), corresponding to the first-order partial derivatives of the value function with respect to the state variables.
@@
Modern machine learning frameworks such as PyTorch allow these costates and their derivatives—including second-order partial derivatives of the value function—to be computed automatically via BPTT and automatic differentiation. We show that a single forward simulation, combined with BPTT, yields unbiased estimates of the costate processes required by PMP. As a result, PG-DPO avoids explicit solution of the backward equation, thereby sidestepping the major bottlenecks of dynamic programming. Both the costates and the value function emerge naturally as byproducts of the learning process.
A central innovation is the Projected PG-DPO (P-PGDPO) variant. P-PGDPO builds on PG-DPO by substituting the estimated derivatives of the value function directly into the structural first-order conditions (FOC) of PMP. During training, costate estimates typically stabilize well before the policy networks are fully converged. P-PGDPO leverages this property by projecting the stabilized costates and their derivatives onto the analytical manifold defined by PMP’s first-order conditions.
The algorithm proceeds in two stages. In Stage 1, the networks are trained until the costate estimates and their derivatives stabilize. In Stage 2, these estimates are substituted into the first-order conditions to construct the optimal policies. The resulting controls adhere closely to the structural form of the true solution, as dictated by PMP, and are recovered quasi-analytically at each time–state point. This requires minimal additional training while substantially reducing computational cost. By disentangling costate estimation from policy construction, P-PGDPO achieves significant gains in both accuracy and scalability.
@@
We provide a theoretical justification for the P-PGDPO method in Theorem 4.2. When the objective function surface is relatively flat near the optimum, the first-order condition (FOC) violation for the policy obtained in the warm-up stage tends to be small. We can show that if this FOC deviation is small and the costate estimates obtained via BPTT are sufficiently accurate, then the control policy recovered in the second stage is close to the true optimal policy.
We demonstrate the numerical accuracy of our methods by comparing them to analytic benchmarks under affine dynamics, where the true solution is available from previous work (Kim and Omberg, , 1996; Liu, , 2007). We further compare against the Deep BSDE approach, which is representative of existing techniques for solving high-dimensional HJB equations. Finally, we apply P-PGDPO to long-horizon investment problems and to models with non-affine return dynamics, thereby demonstrating its broader applicability.
Our numerical experiments in Section 5 show that baseline PG-DPO produces solutions comparable to those of Deep BSDE in terms of mean squared error. However, the P-PGDPO variant yields significantly smaller errors. This improvement arises because PG-DPO, like other policy optimization algorithms, struggles to reliably discover the true optimum, whereas P-PGDPO overcomes this limitation by directly enforcing the structural first-order conditions of PMP. As a result, P-PGDPO not only scales to problems with many assets and state variables, but also accurately disentangles and reconstructs complex intertemporal hedging demands. Furthermore, it generates stable solutions in non-affine settings and converges smoothly to the affine benchmark as the nonlinear terms vanish. By contrast, the Deep BSDE approach—highly successful in other high-dimensional applications such as option pricing and hedging—fails to capture the full structure of the optimal policy, particularly the intertemporal hedging component.
Conceptually, our approach is related to recent methods that parameterize control policies with neural networks and optimize the expected objective function directly through simulation and gradient ascent, thereby bypassing the need to solve the HJB equation (Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024). Our PMP-based PG-DPO framework, however, is explicitly guided by the continuous-time optimality conditions of PMP. Unlike standard reinforcement learning algorithms such as Proximal Policy Optimization (PPO) or Trust Region Policy Optimization (TRPO), which rely on sampled policy gradient estimates, our method computes exact gradients via BPTT. This yields lower-variance updates and improved computational efficiency in high dimensions (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
This work also contributes to the literature on dynamic portfolio choice, which seeks analytical or numerical solutions. Most contributions adopt a discrete-time formulation, with the exception of affine continuous-time models. For example, Campbell and Viceira, (1999, 2001) employ log-linear approximations, while Balduzzi and Lynch, (1999), Lynch and Balduzzi, (2000), and Lynch, (2001) use discrete-space approximations with transaction costs. Garlappi and Skoulakis, (2010) apply state-variable decomposition, and Brandt et al., (2005) propose a Monte Carlo projection of simulated marginal utility onto basis functions—an approach reminiscent of our forward-simulation step, though lacking the scalability and structural guidance of PMP.
In continuous time, semi-analytical solutions are available under affine dynamics, where the HJB reduces to ODEs (Kim and Omberg, , 1996; Liu, , 2007; Buraschi et al., , 2010). For non-affine dynamics, no general solution exists.
@@
Parallel efforts apply reinforcement learning to continuous-time portfolio problems, including policy evaluation, policy gradients, $Q$-learning, and PPO/TRPO methods (Wang et al., , 2020; Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou, , 2023; Zhao et al., , 2023). These have been applied to dynamic mean-variance analysis (Wang and Zhou, , 2020; Dai et al., , 2023; Huang et al., , 2024) and to Merton-type portfolio problems (Dai et al., , 2025). Such approaches learn return dynamics from data, but are typically restricted to few assets and factors. For example, Geng et al., (2023) propose the FaLPO method, which combines policy learning with model calibration but remains limited in scale. In a related line, Cong et al., (2022) introduce the AlphaPortfolio framework, which bypass preliminary estimation of return distributions and directly constructs portfolios using deep reinforcement learning, showing strong out-of-sample performance. By contrast, we assume the data-generating process is known and focus on solving the general continuous-time consumption and portfolio choice problem with many risky assets and multiple state variables. In a complementary direction, Doh et al., (2025) show that reinforcement learning can recover the rational expectations equilibrium in information-based asset market models without relying on explicit market data estimation, providing a theoretical foundation for the use of reinforcement learning in equilibrium settings.
Other contributions closer to ours include Davey and Zheng, (2022), who develop a BSDE-based method with portfolio constraints; Duarte et al., (2024), who propose a machine learning approach for infinite-horizon settings; and Cheridito et al., (2025), who combine generalized policy iteration with PINNs for finite-horizon DP. These methods, however, either require explicit BSDE solutions, restrict attention to stationary infinite-horizon problems, or fail to recover the structural decomposition of myopic and intertemporal hedging demands in high dimensions. Our approach, by contrast, recovers partial derivatives directly through automatic differentiation of costates, enabling accurate estimation of both the myopic and hedging components. [^2] As Cochrane emphasizes, partial derivatives of the value function are critical for long-term portfolio choice yet notoriously difficult to obtain: “Merton’s theory is also devilishly hard to implement, which surely helps to account for its disuse in practice… What are the partial derivatives of the investor’s value function that guide state-variable investments? That is harder still…” (Cochrane, , 2022, pp. 4–5). The aforementioned literature does not demonstrate accurate recovery of intertemporal hedging terms when both the number of assets and state variables are large. [^3] To our knowledge, the existing literature has demonstrated the ability to solve problems involving at most a dozen risky assets and a dozen state variables, with two notable exceptions. Davey and Zheng, (2022) consider a setting with 20 assets, but under a constant investment opportunity set, i.e., without state variables influencing asset returns. Cheridito et al., (2025) address a case with 25 assets, although without providing evidence regarding the precision of the resulting solution. In contrast, our method substantially improves scalability, handling much larger values of $n$ and comparably high values of $k$, while preserving accuracy even in high-dimensional settings. Our companion paper demonstrates that our method is applicable to constrained portfolio optimization problems involving up to 1000 assets (Huh et al., , 2025).